OpenStudy (anonymous):

@AnswerMyQuestions... And @ro561man 5 1/2 /5 1/4? Please help..

OpenStudy (andrewthecookie):

theres no sign like + or -?

OpenStudy (andrewthecookie):

is that the whole question?

OpenStudy (anonymous):

Yeah.. Let me see if I can make it clearer..

OpenStudy (anonymous):

Dividing?

OpenStudy (igreen):

@AndrewTheCookie I believe the question is \(5\dfrac{1}{2} \div 5 \dfrac{1}{4}\)

OpenStudy (andrewthecookie):

Ok @AnswerMyQuestions Thx

OpenStudy (anonymous):

:p

OpenStudy (andrewthecookie):

>:p

OpenStudy (anonymous):

Yeah the / is dividing..

OpenStudy (andrewthecookie):

1. Can the fractions into improper fractions

OpenStudy (andrewthecookie):

Change* sorry

OpenStudy (anonymous):

That's alright... :)

OpenStudy (andrewthecookie):

Can you do that yourself?

OpenStudy (andrewthecookie):

I just saw the iGreen gave me that answer not Answer xD

OpenStudy (andrewthecookie):

Tootles you there?

OpenStudy (anonymous):

Um... Well would the answer be 2/1 and 4/1? Yes I am here..

OpenStudy (andrewthecookie):

No lemme explain:

OpenStudy (andrewthecookie):

5 1/2 would equal 11/2

OpenStudy (andrewthecookie):

and 5 1/4 = 16/4

OpenStudy (andrewthecookie):

so now it equals 11/2 ÷ 16/4

OpenStudy (andrewthecookie):

to do these problem you must flip the second value (16/4) to 4/16 and then multiply straight across

OpenStudy (andrewthecookie):

now it equals 11/2 x 4/16

OpenStudy (andrewthecookie):

11 x 4 = 44

OpenStudy (andrewthecookie):

2 x 16 = 32

OpenStudy (andrewthecookie):

the answer is 44/32

OpenStudy (andrewthecookie):

you may simplfy if needed

OpenStudy (anonymous):

Umm.. Okay? I understand a little..

OpenStudy (andrewthecookie):

ok good :D

OpenStudy (andrewthecookie):

if u need more help tag me

OpenStudy (anonymous):

But I do... Lol.. Is it because you want me to do it in a another open question? Cause I can..

OpenStudy (andrewthecookie):

ok

OpenStudy (anonymous):

Okay what?

OpenStudy (andrewthecookie):

post the Q

OpenStudy (anonymous):

An open one? Like a new one?

OpenStudy (andrewthecookie):

well ima go so can we do this later? just fan me or pm me

OpenStudy (anonymous):

Okay...? Umm.. I'll give you a medal.. Thanks for your help and time...

OpenStudy (andrewthecookie):

ok